# Intents, Techniques, and Components: a Unified Analysis of Interaction Authoring Tasks in Data Visualization

### v2

@@ 5.1.2. Inconsistencies within the Same Taxonomy
We also found that the categories from the same taxonomy can focus on different aspects of interactivity. Consider the two examples in Figure 3. Figure 3(a) (Liu et al., 2013) shows two bar charts, brushing over a month bar in the left chart will show the daily distribution of that month in orange in the right chart. Figure 3(b) (The Economist, [n. d.]) shows a list of currencies in the left table, with a base currency (US dollar) colored in dark gray. The lollipop chart on the right shows all the currencies’ Big Mac index against the base currency. Selecting a currency (Swiss franc) in the table will highlight the corresponding lollipop. Using Yi et al.’s taxonomy, we can label the interactions in both examples as connect: to “highlight associations and relationships between data items that are already represented” in different views (Yi et al., 2007). However, if we focus on how the targets are changing, the two examples can be categorized differently. In Figure 3(a), the target (orange histograms) updates to show a subset of the data matching a specific month, hence it can be categorized as filter: “change the set of data items being presented based on some specific conditions” (Yi et al., 2007); in Figure 3(b), on the other hand, the target (lollipop) is just a different visual representation of the same data item (Switzerland Franc), and it is differentiated from the other lollipops using fill color, hence matching the definition of select: “by making items of interest visually distinctive, users can easily keep track of them even in a large data set and/or with changes in representations” (Yi et al., 2007).
@@ 6. Technique Categories
We consolidate techniques from existing taxonomies and incorporate additional techniques from our example analysis. Our primary motivation is to reference taxonomies that focus on target change-centric perspective when defining intents. Thus, we also began the amalgamation process using Yi’s and Heer and Shneiderman’s taxonomies (Yi et al., 2007; Heer and Shneiderman, 2012). Our initial step was to gather techniques associated with user intents from each paper. The Table 1 shows the collection. In consolidating these techniques, we further segment them into more granular levels. For example, while both papers define Select as a user intent, they do not delve into the detailed techniques. To address this, we include the related techniques with the user intent, Select (i.e., Point Select, Multi-Select, Range Select, Generalized Select, Linked Select, and Deselect), all informed by our example analysis. More techniques can be added as needed; thus, we define a group of primary techniques in Section 8, presenting a unified analysis.
Yi et al. (Yi et al., 2007) Heer and Shneiderman (Heer and Shneiderman, 2012)
@@ 7. Component Categories
In this section, we present the details on our analysis at the component level. The focus of our analysis at this level is: given a triggering mouse or keyboard event and a target, to attain the observed visual changes in the target, what visualization components are involved, and how are the components used or updated? We first review recent literature that examines the design space of static data visualizations (Zhicheng Liu et al., 2021; Satyanarayan et al., 2019; Liu et al., 2018; Ren et al., 2018; Thompson et al., 2020) and synthesize the components identified in these papers. Given the static components, we then used Vega-Lite’s interaction grammar to guide our top-down analysis of interactive behavior. In particular, we found the selection primitives to be powerful concepts to describe interaction behavior. Our analysis adapted and extends these concepts in two ways. First, we try to identify the components of interactivity with a focus on the input and output of each component. Second, the Vega-Lite does not provide an in-depth account of the role of UI controls, which are commonly used in real-world examples (26 out of 47 in our collection), and we introduced field and component references to address this issue. Below we introduce the component types uncovered from our analysis. We will elaborate on how the manipulation of these components relate to higher-level techniques and intents in Section 8.
@@ 7.2. Observable Interaction Components
Hit Object. Given a mouse event received by a canvas listener, most visualizations conduct a hit-test (Foley et al., 1996) to determine if any visual object intersects with mouse cursor. The result of the test, if not null, is a hit object. A hit object is usually a mark or a glyph.
@@ 7.3. Internal Components
Camera. A component providing a vantage point of the visual representation. A camera has a number of configurable attributes, such as focus point (center of camera), field of view (extent visible through the camera), zoom level, and rotation. In web-based 2D visualizations, the camera is often implemented as the SVG view box (MDN Web Docs, [n. d.]) that can pan and zoom; in 3D visualizations, perspective cameras are used widely (ScienceDirect Engineering Topics, [n. d.]).
@@ 8.1.1. Select: Visually Differentiate Objects from Surroundings
In point selection, a visual object is marked and visually differentiated from its surroundings. For example, in Figure LABEL:fig:pointselect (Alexander, nd), clicking on a glyph representing the date 12/21/2017 changes its stroke color and width, and makes the other glyphs opaque. To support point select, evaluate every target visual object to determine if it is the hit object; alternatively, create a predicate (e.g., data == 12/21/2017), and evaluate the targets against the predicate. The evaluation result is then used to encode the visual channels. Multi-select works in a similar way, except that the predicate operand is a list of data values, and the operator is in. For example, in Figure LABEL:fig:multiselect (Parlapiano and Ashkenas, 2014), six selected polylines representing six industries are highlighted in green, while the other polylines turn gray and opaque. Point select and multi-select can be triggered by mouse events such as clicking (Figure LABEL:fig:pointselect) and scrolling (Figure LABEL:fig:multiselect), as well as UI events such as a button click.
@@
To range select visual objects along the x or y dimension (e.g., Figure 4(a)) (Vega-Lite, [n. d.]), either (1) evaluate every target visual object to determine if it intersects with pixel range defined by the mouse event (click+drag); or (2) identify the data field encoding the position (e.g., date in Figure 4(a)), convert the pixel range into a predicate (e.g., date $\in$ [3/20, 8/6]), and evaluate if each target’s date value satisfies the predicate. The evaluation result is used to encode the visual channel values, where targets failing the evaluation turn gray.
@@
Generalized selection (Heer et al., 2008) expands the hit object to a broader set of visual objects. For example, hovering over an arc representing a webpage click selects the arc as well as arcs representing the previous clicks in a sunburst chart (Figure 4(b)) (Kerry, [n. d.]). To support generalized selection, create a predicate based on the hit object (e.g., clickID $\in$ [sequence from root to “c5”]), then evaluate each target against the predicate, finally, encode the visual channel (i.e., opacity) values according to the evaluation result.
@@ 8.1.2. Annotate: Show Auxiliary Visual Objects
A reference line is often shown to indicate the predicate value used for selection, or to facilitate visual estimation of data values. For example, in Figure 5(a) (Wolf, [n. d.]), hovering on a set of related charts sharing the same x-axis shows a vertical line to help users keep track of the location of their action and the currently selected date. In another example (Figure 5(b)) (Scatterplot with Voronoi, [n. d.]), hovering over a circle shows two orthogonal orange lines, intended to help users read the x and y values more accurately. In our analysis, we assume that these reference lines have been created, and interaction primarily changes their visual channels such as visibility and position. To enable updating reference object, different components are involved depending on the visual channels and chart design. For example, in Figure 5(a), the vertical line’s position is based on mouse x-coordinate, and its visibility depends on whether a valid date predicate operand is found based on mouse x-coordinate; in Figure 5(b), the lines’ positions are based on mouse coordinates, and their visibility depends on whether a hit object is found.
@@ 8.2.1. Steer: Direct Camera View
Geometric zoom updates the visible area of a scene by reconfiguring camera focus point and zoom level based on mouse events. Figure 6 (Bostock, [n. d.]) shows a zoomable bar chart. Users double click anywhere in the chart or pinch through mouse wheel to perform zooming. To enable geometric zoom, the coordinates of the input mouse event typically become the new focus point. The zoom level is usually updated by fixed increments based on the amount of pinching or the number of double clicking.
@@
([Uncaptioned image][Uncaptioned image]) In tab-based user interfaces, it is common to display a view at a time by clicking on the UI control. For example, in GOV — DNA, by clicking on one of the four buttons at the top (e.g., Figure 8(a)), the currently visible view is updated. The views are usually predefined, each containing unique definitions of components, such as marks, encodings, and even predicates (to control which data items are visualized). Depending on how the views are defined, this technique can be realized by updating either a state variable or the camera. In Figure 8(a), the views are significantly different, and it is cleaner to treat them as independent scene definitions. Toggling between them thus requires the definition of a state variable currentScene. In examples such as the Star Mapper (Interactive, [n. d.]), the toggling is more easily described by updating the camera’s position and focus point in a three-dimensional scene.
@@ 8.3.1. Encode: Change the Visual Representation of Existing Data Items
In Figure 8(b) (Laura, [n. d.]), clicking on the buttons (“Total” or “Compare”) toggles the visualization between a multi-line graph and a stacked area chart, showing the same data items. Multiple aspects of visual encoding are changed: mark type, visual channels and layout. To implement changing chart type, define state variables referencing the chart type definition and associated encodings, and create encoders that handle the mapping between data values to the corresponding marks and channels.
@@ 8.3.2. Enter/Exit
In our example collection, we only encountered one case that introduces new data items through interaction. In a D3.js demo (Liu, 2004), clicking the background will add data points which are shown as new circle marks at the clicked locations. This example shows how interaction can modify the data being visualized.
@@ 8.4.1. Filter
Dynamic Queries (Ahlberg et al., 1992; Shneiderman, 1994) allow users to control which data items should be visualized through interaction with user interface elements such as sliders, checkboxes, and manipulable widgets (Hochheiser and Shneiderman, 2004; Wattenberg, 2001) that embody filtering criteria. To support dynamic queries, define a predicate for each querying control, and update predicate operands based on user interaction. The target visual objects can then be updated in two ways. First, we can evaluate the targets against a predicate evaluator, and encode the targets’ visibility accordingly, but this approach will not work for visualizations that require re-computation of layouts (e.g., stacked area chart in Figure 8(b)). Second, we can treat the scene as the target and update its data based on the predicate. The scene can then be updated with the same encoding mechanisms.
@@ 8.4.2. Abstract/Elaborate: Change the Levels of Abstraction
Given a hierarchical dataset, users can choose whether to display the entire structure (including all the leaf nodes and nodes at the intermediate levels), or only nodes above a certain level. For example, DendroMap (Bertucci et al., 2022) displays hierarchical clusters of image datasets. Users can increase or decrease the number of visible clusters (Figure 9) through a slider. The interaction sets the level of abstraction in the underlying hierarchy. To enable moving up and down a hierarchy, create a state variable recording the desired level of abstraction, and update the target data whenever the variable value changes.
@@
Levels of abstraction can be changed on multivariate dataset too. In Polestar (Moritz et al., [n. d.]), Polaris (Stolte et al., 2002) and Tableau (Tableau Software, [n. d.]), users can drill down data by dragging additional dimensions to the row or column shelf, which will partition existing data displays into a finer grained aggregation. To enable drill-down and roll-up, create state variables recording the current dimensions and measure, and update the target data (based on pre-computed data cubes or real-time querying) whenever the variable values change.
@@ 8.4.3. Derive: Compute and Show New Field Values from Data Items Attached to Existing Objects
In visualizations where a mark represents multiple data items, it is possible to change how the values for a field are aggregated. For example, Polestar (Moritz et al., [n. d.]), Polaris (Stolte et al., 2002) and Tableau (Tableau Software, [n. d.]) allow users to choose from a list of aggregation functions (Figure 10). To support changing aggregator, define a state variable representing the aggregator, and update the target scene’s data whenever the aggregator is changed.
@@ 8.5. Multiple User Intents
Unlike geometric zoom, semantic zoom has dual intents: steer and abstract/elaborate. When zooming in the heatmap in Figure 11(a) (Heavy.ai, [n. d.]), not only the camera zoom level changes, the bin size of the heatmap also updates to show more granular information.
@@
Direct walk “smoothly move[s] the viewing focus from one position in information structure to another” (Yi et al., 2007). In Figure 11(b) (Kerr, [n. d.]), clicking on a table row both selects a county map polygon and performs a geometric zoom with the county as the focus point.
@@ 9. Discussion: Descriptive, Evaluative, and Generative Powers and Completeness
We evaluate the descriptive, evaluative, and generative powers (Beaudouin-Lafon, 2004) of our framework, and discuss the completeness of the taxonomies.
@@ 9.1. Descriptive Power
To understand how well our framework can describe complex and bespoke interaction designs, we choose Dust and Magnet (DnM) (Soo Yi et al., 2005) and OnSet (Sadana et al., 2014) as two case studies, based on the following considerations: (1) they use customized static visual representations instead of standard chart types, and (2) the interactive behaviors deviate from the representative techniques included in our analysis.
@@ 9.2. Evaluative Power
Levels of abstraction. The first aspect, levels of abstraction is to examine where a system primarily operates at. Each system encompasses different levels of abstraction, which defines the range of interaction authoring in the system. For example, Interactive+ is to augment an existing static web-based visualization with interactivity. This tool operates primarily at the component level: given a visualization realized in the HTML Document Object Model (WHATWG, nd), Interaction+ extracts component-level information including the visual objects and their attributes as well as the distribution of visual objects over each visual attribute. Interaction+’s user interface does not provide explicit abstractions of concepts at the technique and user intent level. As an example of a system that covers multiple abstraction levels, consider Lyra 2, which grants access to two levels of abstraction. At the component level, author’s demonstration is primarily in the form of choosing visual marks in a static visualization generated by Vega or Vega-Lite, and constructing query objects that are translated to predicates. An interaction inspector also provides component-level information about recommended design (e.g., mouse event, visual attributes); at the technique level, the suggested interaction techniques are shown as selectable widgets with visual thumbnails in the inspector. Lyra 2 does not provide explicit visual representations or interaction mechanisms at the user intent level.

### v3

@@ 5.1.2. Inconsistencies within the Same Taxonomy
We also found that the categories from the same taxonomy can focus on different aspects of interactivity. Consider the two examples in Figure 3. Figure 3(a) (liu_immens_2013) shows two bar charts, brushing over a month bar in the left chart will show the daily distribution of that month in orange in the right chart. Figure 3(b) (the_economist_big_nodate) shows a list of currencies in the left table, with a base currency (US dollar) colored in dark gray. The lollipop chart on the right shows all the currencies’ Big Mac index against the base currency. Selecting a currency (Swiss franc) in the table will highlight the corresponding lollipop. Using Yi et al.’s taxonomy, we can label the interactions in both examples as connect: to “highlight associations and relationships between data items that are already represented” in different views (yi_toward_2007). However, if we focus on how the targets are changing, the two examples can be categorized differently. In Figure 3(a), the target (orange histograms) updates to show a subset of the data matching a specific month, hence it can be categorized as filter: “change the set of data items being presented based on some specific conditions” (yi_toward_2007); in Figure 3(b), on the other hand, the target (lollipop) is just a different visual representation of the same data item (Switzerland Franc), and it is differentiated from the other lollipops using fill color, hence matching the definition of select: “by making items of interest visually distinctive, users can easily keep track of them even in a large data set and/or with changes in representations” (yi_toward_2007).
@@ 6. Technique Categories
We consolidate techniques from existing taxonomies and incorporate additional techniques from our example analysis. Our primary motivation is to reference taxonomies that focus on target change-centric perspective when defining intents. Thus, we also began the amalgamation process using Yi’s and Heer and Shneiderman’s taxonomies (yi_toward_2007; heer_interactive_2012). Our initial step was to gather techniques associated with user intents from each paper. The Table 1 shows the collection. In consolidating these techniques, we further segment them into more granular levels. For example, while both papers define Select as a user intent, they do not delve into the detailed techniques. To address this, we include the related techniques with the user intent, Select (i.e., Point Select, Multi-Select, Range Select, Generalized Select, Linked Select, and Deselect), all informed by our example analysis. More techniques can be added as needed; thus, we define a group of primary techniques in Section 8, presenting a unified analysis.
Yi et al. (yi_toward_2007) Heer and Shneiderman (heer_interactive_2012)
@@ 7. Component Categories
In this section, we present the details on our analysis at the component level. The focus of our analysis at this level is: given a triggering mouse or keyboard event and a target, to attain the observed visual changes in the target, what visualization components are involved, and how are the components used or updated? We first review recent literature that examines the design space of static data visualizations (zhicheng_liu_atlas_2021; satyanarayan_critical_2019; liu_data_2018; ren_charticulator_2018; thompson_understanding_2020) and synthesize the components identified in these papers. Given the static components, we then used Vega-Lite’s interaction grammar to guide our top-down analysis of interactive behavior. In particular, we found the selection primitives to be powerful concepts to describe interaction behavior. Our analysis adapted and extends these concepts in two ways. First, we try to identify the components of interactivity with a focus on the input and output of each component. Second, the Vega-Lite does not provide an in-depth account of the role of UI controls, which are commonly used in real-world examples (26 out of 47 in our collection), and we introduced field and component references to address this issue. Below we introduce the component types uncovered from our analysis. We will elaborate on how the manipulation of these components relate to higher-level techniques and intents in Section 8.
@@ 7.2. Observable Interaction Components
Hit Object. Given a mouse event received by a canvas listener, most visualizations conduct a hit-test (foley_computer_1996) to determine if any visual object intersects with mouse cursor. The result of the test, if not null, is a hit object. A hit object is usually a mark or a glyph.
@@ 7.3. Internal Components
Camera. A component providing a vantage point of the visual representation. A camera has a number of configurable attributes, such as focus point (center of camera), field of view (extent visible through the camera), zoom level, and rotation. In web-based 2D visualizations, the camera is often implemented as the SVG view box (mdn_web_docs_viewbox_nodate) that can pan and zoom; in 3D visualizations, perspective cameras are used widely (sciencedirect_engineering_topics_perspective_nodate).
@@ 8.1.1. Select: Visually Differentiate Objects from Surroundings
In point selection, a visual object is marked and visually differentiated from its surroundings. For example, in Figure LABEL:fig:pointselect (alexander_bank_nodate), clicking on a glyph representing the date 12/21/2017 changes its stroke color and width, and makes the other glyphs opaque. To support point select, evaluate every target visual object to determine if it is the hit object; alternatively, create a predicate (e.g., data == 12/21/2017), and evaluate the targets against the predicate. The evaluation result is then used to encode the visual channels. Multi-select works in a similar way, except that the predicate operand is a list of data values, and the operator is in. For example, in Figure LABEL:fig:multiselect (parlapiano_how_2014), six selected polylines representing six industries are highlighted in green, while the other polylines turn gray and opaque. Point select and multi-select can be triggered by mouse events such as clicking (Figure LABEL:fig:pointselect) and scrolling (Figure LABEL:fig:multiselect), as well as UI events such as a button click.
@@
To range select visual objects along the x or y dimension (e.g., Figure 4(a)) (vega-lite_seattle_nodate), either (1) evaluate every target visual object to determine if it intersects with pixel range defined by the mouse event (click+drag); or (2) identify the data field encoding the position (e.g., date in Figure 4(a)), convert the pixel range into a predicate (e.g., date $\in$ [3/20, 8/6]), and evaluate if each target’s date value satisfies the predicate. The evaluation result is used to encode the visual channel values, where targets failing the evaluation turn gray.
@@
Generalized selection (heer_generalized_2008) expands the hit object to a broader set of visual objects. For example, hovering over an arc representing a webpage click selects the arc as well as arcs representing the previous clicks in a sunburst chart (Figure 4(b)) (kerry_sequences_nodate). To support generalized selection, create a predicate based on the hit object (e.g., clickID $\in$ [sequence from root to “c5”]), then evaluate each target against the predicate, finally, encode the visual channel (i.e., opacity) values according to the evaluation result.
@@ 8.1.2. Annotate: Show Auxiliary Visual Objects
A reference line is often shown to indicate the predicate value used for selection, or to facilitate visual estimation of data values. For example, in Figure 5(a) (wolf_alpha_nodate), hovering on a set of related charts sharing the same x-axis shows a vertical line to help users keep track of the location of their action and the currently selected date. In another example (Figure 5(b)) (scatterplot_with_voronoi_life_nodate), hovering over a circle shows two orthogonal orange lines, intended to help users read the x and y values more accurately. In our analysis, we assume that these reference lines have been created, and interaction primarily changes their visual channels such as visibility and position. To enable updating reference object, different components are involved depending on the visual channels and chart design. For example, in Figure 5(a), the vertical line’s position is based on mouse x-coordinate, and its visibility depends on whether a valid date predicate operand is found based on mouse x-coordinate; in Figure 5(b), the lines’ positions are based on mouse coordinates, and their visibility depends on whether a hit object is found.
@@ 8.2.1. Steer: Direct Camera View
Geometric zoom updates the visible area of a scene by reconfiguring camera focus point and zoom level based on mouse events. Figure 6 (bostock_zoomable_nodate) shows a zoomable bar chart. Users double click anywhere in the chart or pinch through mouse wheel to perform zooming. To enable geometric zoom, the coordinates of the input mouse event typically become the new focus point. The zoom level is usually updated by fixed increments based on the amount of pinching or the number of double clicking.
@@
([Uncaptioned image][Uncaptioned image]) In tab-based user interfaces, it is common to display a view at a time by clicking on the UI control. For example, in GOV — DNA, by clicking on one of the four buttons at the top (e.g., Figure 8(a)), the currently visible view is updated. The views are usually predefined, each containing unique definitions of components, such as marks, encodings, and even predicates (to control which data items are visualized). Depending on how the views are defined, this technique can be realized by updating either a state variable or the camera. In Figure 8(a), the views are significantly different, and it is cleaner to treat them as independent scene definitions. Toggling between them thus requires the definition of a state variable currentScene. In examples such as the Star Mapper (interactive_esa_nodate), the toggling is more easily described by updating the camera’s position and focus point in a three-dimensional scene.
@@ 8.3.1. Encode: Change the Visual Representation of Existing Data Items
In Figure 8(b) (laura_namegrapher_nodate), clicking on the buttons (“Total” or “Compare”) toggles the visualization between a multi-line graph and a stacked area chart, showing the same data items. Multiple aspects of visual encoding are changed: mark type, visual channels and layout. To implement changing chart type, define state variables referencing the chart type definition and associated encodings, and create encoders that handle the mapping between data values to the corresponding marks and channels.
@@ 8.3.2. Enter/Exit
In our example collection, we only encountered one case that introduces new data items through interaction. In a D3.js demo (mouseEvent_Will), clicking the background will add data points which are shown as new circle marks at the clicked locations. This example shows how interaction can modify the data being visualized.
@@ 8.4.1. Filter
Dynamic Queries (ahlberg_dynamic_1992; shneiderman_dynamic_1994) allow users to control which data items should be visualized through interaction with user interface elements such as sliders, checkboxes, and manipulable widgets (hochheiser_dynamic_2004; wattenberg_sketching_2001) that embody filtering criteria. To support dynamic queries, define a predicate for each querying control, and update predicate operands based on user interaction. The target visual objects can then be updated in two ways. First, we can evaluate the targets against a predicate evaluator, and encode the targets’ visibility accordingly, but this approach will not work for visualizations that require re-computation of layouts (e.g., stacked area chart in Figure 8(b)). Second, we can treat the scene as the target and update its data based on the predicate. The scene can then be updated with the same encoding mechanisms.
@@ 8.4.2. Abstract/Elaborate: Change the Levels of Abstraction
Given a hierarchical dataset, users can choose whether to display the entire structure (including all the leaf nodes and nodes at the intermediate levels), or only nodes above a certain level. For example, DendroMap (bertucci_dendromap_2022) displays hierarchical clusters of image datasets. Users can increase or decrease the number of visible clusters (Figure 9) through a slider. The interaction sets the level of abstraction in the underlying hierarchy. To enable moving up and down a hierarchy, create a state variable recording the desired level of abstraction, and update the target data whenever the variable value changes.
@@
Levels of abstraction can be changed on multivariate dataset too. In Polestar (moritz_polestar_nodate), Polaris (stolte_polaris_2002) and Tableau (tableau_software_tableau_nodate), users can drill down data by dragging additional dimensions to the row or column shelf, which will partition existing data displays into a finer grained aggregation. To enable drill-down and roll-up, create state variables recording the current dimensions and measure, and update the target data (based on pre-computed data cubes or real-time querying) whenever the variable values change.
@@ 8.4.3. Derive: Compute and Show New Field Values from Data Items Attached to Existing Objects
In visualizations where a mark represents multiple data items, it is possible to change how the values for a field are aggregated. For example, Polestar (moritz_polestar_nodate), Polaris (stolte_polaris_2002) and Tableau (tableau_software_tableau_nodate) allow users to choose from a list of aggregation functions (Figure 10). To support changing aggregator, define a state variable representing the aggregator, and update the target scene’s data whenever the aggregator is changed.
@@ 8.5. Multiple User Intents
Unlike geometric zoom, semantic zoom has dual intents: steer and abstract/elaborate. When zooming in the heatmap in Figure 11(a) (heavyai_1990_nodate), not only the camera zoom level changes, the bin size of the heatmap also updates to show more granular information.
@@
Direct walk “smoothly move[s] the viewing focus from one position in information structure to another” (yi_toward_2007). In Figure 11(b) (kerr_d3_nodate), clicking on a table row both selects a county map polygon and performs a geometric zoom with the county as the focus point.
@@ 9. Discussion: Descriptive, Evaluative, and Generative Powers and Completeness
We evaluate the descriptive, evaluative, and generative powers (beaudouin-lafon_designing_2004) of our framework, and discuss the completeness of the taxonomies.
@@ 9.1. Descriptive Power
To understand how well our framework can describe complex and bespoke interaction designs, we choose Dust and Magnet (DnM) (soo_yi_dust_2005) and OnSet (sadana_onset_2014) as two case studies, based on the following considerations: (1) they use customized static visual representations instead of standard chart types, and (2) the interactive behaviors deviate from the representative techniques included in our analysis.
@@ 9.2. Evaluative Power
Levels of abstraction. The first aspect, levels of abstraction is to examine where a system primarily operates at. Each system encompasses different levels of abstraction, which defines the range of interaction authoring in the system. For example, Interactive+ is to augment an existing static web-based visualization with interactivity. This tool operates primarily at the component level: given a visualization realized in the HTML Document Object Model (whatwg_dom_nodate), Interaction+ extracts component-level information including the visual objects and their attributes as well as the distribution of visual objects over each visual attribute. Interaction+’s user interface does not provide explicit abstractions of concepts at the technique and user intent level. As an example of a system that covers multiple abstraction levels, consider Lyra 2, which grants access to two levels of abstraction. At the component level, author’s demonstration is primarily in the form of choosing visual marks in a static visualization generated by Vega or Vega-Lite, and constructing query objects that are translated to predicates. An interaction inspector also provides component-level information about recommended design (e.g., mouse event, visual attributes); at the technique level, the suggested interaction techniques are shown as selectable widgets with visual thumbnails in the inspector. Lyra 2 does not provide explicit visual representations or interaction mechanisms at the user intent level.